OpenStudy (anonymous):

Simplify: 4x − 3y + 2z − 3x − y + 3z = ?

OpenStudy (anonymous):

Combine like terms.

OpenStudy (anonymous):

Combine your x-values together- 4x-3x=x Combine your y-values together- -3y-y=-3y+(-y)=-4y Combine your z-values together- 2z+3z=5z This gives us- x-4y+z

OpenStudy (anonymous):

(4x - 3x) + (-3y - y) + (2z + 5z) = ?

OpenStudy (anonymous):

Just add all of the common terms together. 4x - 3x - 3y - y + 2z + 3z = x - 4y + 5z x - 4y + 5z is your answer.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

np :)